Solid Iodine is added to a strongly basic solution of mercury (II) nitrate. What are the products and how do you solve it?
1
Expert's answer
2012-10-31T08:22:34-0400
If ahighly alkaline solution add-crystal iodine, the following reactions:

3I2 + 6KOH = 5KI + KIO3+ 3H2O

Next KI reacts with nitrate of mercury to form a yellow precipitate:
Hg(NO3)2 + 2KI = HgI2 + 2 KNO3